OpenStudy (anonymous):

Calculus: f(x)=x, g(x)=x^2 =integral from -1 to 1 of f(x)g(x) dx

OpenStudy (anonymous):

There is no way the answer if pi/2 right?

OpenStudy (anonymous):

seems pretty damned unlikely

OpenStudy (anonymous):

Do you mean:\[\int\limits_{-1}^{1}[f(x)g(x)]dx\]Isn't that equal to:\[\int\limits_{-1}^{1} xx^{2}dx = \int\limits_{-1}^{1} x^{3}dx = 0\]? Am I missing something?

OpenStudy (anonymous):

\[\int_{-1}^1x^3dx=0\]

OpenStudy (anonymous):

i got 0

OpenStudy (anonymous):

yea it is odd, must be zero

OpenStudy (anonymous):

okay, cool thanks

OpenStudy (anonymous):

with this new system, i cannot give more than one medal, sorry @bmp. But thanks to you as well

OpenStudy (anonymous):

@LagrangeSon678 no problem, mate. Glad that I could be of any help :-)